OpenStudy (anonymous):

how do i add (5x/10)+(x+1/2x)?????

OpenStudy (tkhunny):

Have you considered a common denominator?

OpenStudy (shamil98):

or multiplying everything by 2 to get rid of the fractions

OpenStudy (tkhunny):

Is it \(\dfrac{1}{2x}\;or\;\dfrac{1}{2}x\) Don't just multiply by 2. That changes the value.

OpenStudy (anonymous):

i forgot to add an x to the 10. would i have to multiply 5x/10 by 2 and x+1/2 by 10???

OpenStudy (anonymous):

yep, same question. is it \[\frac{ 1 }{ 2 }x OR \frac{ 1 }{ 2x } ? \]

OpenStudy (anonymous):